On January 1st, 2013, Nicklaus Corp issued $5M of 10-year 6% semiannual bonds. The market rate at the
time of the issuance was 4%. The bonds makes interest payments on January 1 and July 1.


1. Did Nicklaus receive more or less cash than the face value of the bonds? By how much?
2. Assuming a December 31st fiscal year end, how much was interest paid in 2013?
3. What will the total interest expense be over the life of the bond?
4. How much was interest expense in 2013 (round to the nearest dollar)?
Answer #1

Solution 1:

Chart Values are based on:
n= (10 Years*2) 20 Half years
i= (4%/2) 2% Semi annual
Cash Flow Table Value * Amount = Present Value
Par (Maturity) Value 0.67297 * $50,00,000 = $33,64,857
Interest (Annuity) [$5,000,000*6%*6/12] 16.35143 * $1,50,000 = $24,52,715
Price of Bonds $58,17,572

Nicklaus receive more cash than the face value of the bonds by = $5817572 - $5000000 = $817,572

Solution 2:

Interest paid in 2013 (at July 1) = $5,000,000*6%*6/12 = $150,000

Solution 3:

Total bond interest expense Over life of bonds:
Amount Paid:
20 Payments of ($5,000,000*6%*6/12=$150,000) each $30,00,000
Par Value at maturity $50,00,000
Total Repaid $80,00,000
Less: Amount Borrowed $58,17,572
Total Bond interest Expense $21,82,428

Solution 4:

First semiannual Interest expense at July 1, 2013 = $5817572 *4% *6/12 = $116,351

First semiannual Premium Amortized = Interest Paid - Interest Expense = $150,000 - $116,351 = $33,649

Second Semiannual Interest expense at 31 Dec, 2013 = ($5817572 - $33649) *4%*6/12 = $115,678

Total interest expense in 2013 = $116,351 + $115,678 = $232,030